Naciama:

Complete the table below to solve the equation 2.5x − 10.5 = 64(0.5x). x f(x) = 2.5x − 10.5 g(x) = 64(0.5x) 2 3 4 5 6 A newspaper started an online version of its paper 14 years ago. In a recent presentation to stockholders, the lead marketing executive states that the revenues for online ads have more than doubled that of the revenues for printed ads since starting the online version of the paper. Use the graph below to justify the lead executive’s statement. Determine the approximate year that the two ad revenues were equal. Two ocean beaches are being affected by erosion. The table shows the width, in feet, of each beach measured at high tide where 1995 is represented by year 0: Year number Western Beach width (in feet) Dunes Beach width (in feet) 0 100 20 5 90 45 10 80 70 11 78 75 12 76 80 15 70 95 Describe the patterns shown by the erosion data measurements shown for each of the beaches in the table. Between which years will the beaches have approximately the same width? Assuming these rates remain constant, what can you do to get a better approximation of when the two beaches will have the same width?

ItzAgar:

Can you clarify what exactly you need help on?

Florisalreadytaken:

hey there! please assure to post a screenshot of your problem as its not very clear, the table. is it 2.5x − 10.5 = 64(0.5x) or 2.5^x − 10.5 = 64(0.5^x) -- there miht be problems with the emponents :)